OpenStudy (ucreategames):

I can't solve this.... Solve for x. 18^(x+8)=13^(-9x)

OpenStudy (anonymous):

hit both sides with "ln"

OpenStudy (anonymous):

if you heard of natural logarithms

OpenStudy (anonymous):

(or log with any other base if you prefer that, base of 10 - which is typical .. or any positive base)

OpenStudy (anonymous):

If you didn't learn about logarithms, you can graph y=18^(x+8) y=13^(-9x) and look for the x value at which the functions will intersect

OpenStudy (ucreategames):

Thank you...

OpenStudy (anonymous):

yw